OpenStudy (anonymous):

m-n/4 if m=5 and n=8

OpenStudy (anonymous):

will the answer be -1

OpenStudy (anonymous):

Put m for 5 and put n for 8 into the equation. (5-8)/4

OpenStudy (anonymous):

ok i know that but do you do PEMDAS after that or no

OpenStudy (anonymous):

What are your answer choices?

OpenStudy (anonymous):

I'm not quite sure if you have to do PEMDAS for this.

OpenStudy (anonymous):

ok and i don't have any answer choices. my teacher just gave us a worksheet to solve the problem

OpenStudy (anonymous):

\[\frac{ (5-8) }{ 4 } = -\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

No problem, and don't forget about the medal. By clicking the best response.